# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ II Floquet prethermalization in the 2D Ising model
The unitary $\mathscr{U}$ of a single Trotter layer can be realized with a quantum circuit consisting of two layers of single-qubit gates enclosing a fixed number of layers of two-qubit gates that depends on the lattice connectivity (Fig. 1b). For the simulations in this work we use a 2D rectangular lattice with periodic boundary conditions, and therefore each Trotter step consists of four layers of ($N/2$) two-qubit gates $U_{ZZ}(2Jdt)$, where $U_{ZZ}(\varphi)=\exp(-i\varphi Z\otimes Z/2)$. The $U_{ZZ}(\varphi)$ gates are implemented natively on the H2 quantum computer (i.e., they are not further decomposed into perfect entanglers).
@@ III Quantum and classical simulations
Quantum and classical simulation results for the low temperature ($\Delta\theta=0$) quench are shown in Fig. 3a. The semi-transparent solid lines are obtained using MPS-based quantum circuit simulations [27] employing a method similar to that described in Ref. [28] (see Sec. A4 for further details). We carried out these simulations up to a maximum MPS bond dimension of $\chi=4000$, for which MPS-based truncation errors were sufficiently small that we could reliably extrapolate the largest-bond dimension results to the limit of zero truncation (purple line, zero-truncation extrapolation or ZTE). We also show raw (completely unprocessed) data from the H2 quantum computer. The dominant source of error in these simulations is imperfections in the two qubit gates, which have an average fidelity of $99.94(1)\%$ (see Sec. A2.2). We also obtained data for noise amplified (NA) circuits, in which two-qubit Pauli operators are stochastically inserted (based on a learned noise channel of our two-qubit gates) in order to artificially reduce the quality of the quantum computation [12]. From the raw and NA data, the quantum data is extrapolated to the limit of zero noise using zero-noise extrapolation (ZNE) [29]. Details of the ZNE methods, along with a novel strategy to remove the impact of detectable leakage errors without post-selection (zero-noise regression, or ZNR), are presented in the appendices. The good agreement of the processed data with the extrapolated MPS results gives us confidence in the error-mitigation strategies in a regime where they have considerable error to remove.
@@
Figure 3c shows the same error-mitigated quantum data for the intermediate temperature quench compared to exact diagonalization results for smaller systems (gray curves), along with further approximate classical methods for the $7\times 8$ lattice, providing a comprehensive comparison of the quantum data with a variety of classical approaches at the frontier of numerical simulations. We explored two different classes of methods: Schrödinger picture methods that evolve a compressed representation of the wave function, and Heisenberg picture methods that evolve a compressed representation of the observable $Z_{\rm tot}^{2}$. For the former, we attempted to evolve the wave function using a projected entangled pair state (PEPS) ansatz and a variety of compression schemes (including full update and belief propagation). We found that evolving a PEPS with periodic-boundary conditions using belief propagation (BP-PEPS), and then computing $\langle Z_{\rm tot}^{2}\rangle$ by sampling, produced accurate observables to the longest times, though we were still not able to obtain converged results beyond about 8 Trotter steps (mainly due to GPU memory limitations, see Sec. A5.2). Simulations in the Heisenberg picture were carried out using both projected entangled pair operator (PEPO, see Sec. A5.3) and sparse-Pauli methods [33, 34, 35, 36, 37, 38] (see Sec. A6 comparing coefficient and Pauli weight truncations). In general, we found PEPO evolution to be less accurate than PEPS simulations with comparable resources. It would be interesting to combine Schrödinger and Heisenberg techniques [33, 39], though the limitations of each independently for this quench suggests that a hybrid approach would likely face challenges in substantially extending the temporal range. Sparse Pauli dynamics simulations retained about 4 billion Pauli operators and used roughly $1{\rm TB}$ of memory, and while they are in principle systematically improvable with more memory they could not be converged with available resources. We also performed neural network quantum state (NQS) simulations using the convolutional neural network (CNN) architecture similar to Ref. [40]. The NQS is time evolved by time-dependent variational Monte Carlo (tVMC) [41] using the NetKet library [42]. By systematically increasing the number of parameters of the ansatz, we obtained good results for smaller systems, but we were not able to obtain quantitatively accurate results except at very short times for a 7x8 lattice (detailed results are presented in Appendix A7). NQS-based methods are a rapidly developing field of study, and it would be interesting to see whether any of the many possible variations can be brought in agreement with the quantum data out to late times.
@@ IV Emergent hydrodynamics
Figure 4: Emergent hydrodynamics. Hydrodynamic relaxation of the spin-exchange contribution to energy density for a quench in an $14\times 4$ strip with an inhomogeneous initial state (described in text). (a) Evolution of the y-averaged exchange energy density profiles, $\mathcal{E}(x)$. (b) Dashed lines show a fit of ~$\tilde{\mathcal{E}}(q)$ to $ae^{-\Gamma_{q}s}+c\delta_{q,0}$. Fits are restricted to long-wavelengths ($q=2\pi n/L_{x}$ for $n=0,1,2,3$), and exclude data points with poor signal-to-shot-noise ratio (unfilled symbols). The extracted decay rates are consistent with the expected diffusive scaling of the heat equation, $\Gamma_{q}\approx\mathcal{D}q^{2}$ (inset) with diffusion constant $\mathcal{D}=0.38(5)$.
An essential aspect of prethermalization in Floquet systems is the emergence of an approximately conserved energy, whose violation only becomes apparent at exponentially late times, $t\gtrsim\tau_{H}$. In chaotic (pre)thermalizing systems, local observables that overlap with the conserved energy density are expected to exhibit a short-time transient relaxation to a local thermal quasi-equilibrium followed by a slow, hydrodynamic relaxation of long-wavelength energy inhomogeneities at late times. To probe the expected emergence of hydrodynamics, we simulate a quench with an inhomogeneous initial condition in which a thin vertical domain of down (|$\ket{1}$) spins is embedded in an otherwise uniform background, corresponding to Eq. (2) with site dependent $\theta_{j}=\pi$ if $x_{j}=L_{x}/2$ or otherwise $\theta_{j}=\theta_{\rm min}+2\pi/9$ ($x_{j}$ being the $x$-coordinate of qubit $j$). To create an appropriate separation of time-scales between the local thermalization time ($t_{0}\gtrsim 5$ time steps for these parameters), and late-time hydrodynamics, we consider an elongated rectangular geometry with dimensions $(L_{x},L_{y})=(14\times 4)$. We measure on the quantum processor the local density of $H_{ZZ}$: $\mathcal{E}_{i}=\frac{J}{4}\sum_{j\in\langle i\rangle}\langle Z_{i}Z_{j}\rangle$, average this quantity over the $y$ coordinate, $\mathcal{E}(x)=\frac{1}{L_{y}}\sum_{i:x_{i}=x}\mathcal{E}_{i}$ (see Fig. 4a), and compute its 1D Fourier transform ~$\tilde{\mathcal{E}}(q)=\sum_{x}\frac{e^{iqx}}{\sqrt{L_{x}}}\mathcal{E}(x)$ (see Fig. 4b). The $q=0$ mode quickly saturates to a steady-state value, indicating its overlap with an emergent conserved energy. The smallest few non-zero ($q_{n}=2\pi n/L_{x}$, $n=1,2,3$) wavevectors relax with characteristic rates $\Gamma_{q}$, consistent with the expected behavior for the heat equation, $\Gamma_{q}\approx\mathcal{D}q^{2}$ (valid asymptotically for long wavelengths, $|{q}|\ll 1$), with diffusion constant $\mathcal{D}=0.38(5)$. Components with larger $q$ decay rapidly within the initial local thermalization and do not survive into the hydrodynamic regime.
@@ Acknowledgments
We acknowledge the entire Quantinuum team for their many contributions towards successful operation of the H2 quantum computer, and Honeywell for fabricating the trap used in this experiment. We thank Konstantinos Meichanetzidis, Chris Langer, Grahame Vittorini, Stephen Erickson, James Hostetter, and David Liefer for helpful comments on the manuscript and thank Dan Mills for useful discussions. Fermioniq was supported by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL Programme. W.K., M.K., and F.P. acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC–2111–390814868, TRR 360 – 492547816 and DFG grants No. KN1254/1-2, KN1254/2-1, FOR 5522 (project-id 499180199), the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 851161), the European Union (grant agreement No 101169765), as well as the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. E.G. acknowledges funding by the Bavarian Ministry of Economic Affairs, Regional Development and Energy (StMWi) under project Bench-QC (DIK0425/01). M.S.R. acknowledges funding from the 2024 Google PhD Fellowship, the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602), and the the Swiss National Science Foundation (grant number 200021-219329). SPD simulation work at Caltech was supported by the US Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. TN-BP algorithm development at Caltech was supported by the US Department of Energy, Office of Science, Basic Energy Sciences and Advanced Scientific Computing Research, Quantum Utility through Advanced Computational Quantum Algorithms.
@@
which performs a rotation by $0\leq\theta\leq\pi$ about a Bloch vector in the $xy$ plane specified by the angle $0\leq\phi<2\pi$. An $X$ gate used in DD can be implemented in two physically inequivalent ways: $U_{1q}(\pi,0)=-iX$ and $U_{1q}(\pi,\pi)=+iX$. While these two gates differ by a logically-irrelevant global phase, coherent errors on the two gates due to systematic under- or over-rotations are not logically equivalent. During DD, we alternate the phase of each $X$ gate on each qubit between $U_{1q}(\pi,0)$ and $U_{1q}(\pi,\pi)$ to approximately cancel such errors.
@@ A Pauli error channel is given by
In this section we present the classical simulation results of the digitized TFIM obtained with 1D matrix-product-state (MPS) [57] methods. The results are accompanied by a quantitative analysis of the extrapolation method used, from which we conclude that, for the observable of interest [$Z_{\rm tot}^{2}$ from Eq. (4)], extrapolating in simulation fidelity is more reliable than extrapolating in bond dimension – an observation that likely holds more generally beyond the scope of this work. Furthermore, based on comparisons to exact simulations for smaller systems, we include a heuristic confidence interval on the extrapolations. The zero-noise extrapolated expectation values obtained from H2 are found to lie within this confidence interval at the early times for which it can be reliably established (i.e., before the MPS fidelity becomes so low that the extrapolations are completely unpredictable, up to $s\approx 8$ for the intermediate quench).
@@
Figure A12: Relative error $\left(\langle Z_{\rm tot}^{2}\rangle-\overline{\langle Z_{\rm tot}^{2}\rangle}\right)/\left|\langle Z_{\rm tot}^{2}\rangle\right|$ between the exact value $\langle Z_{\rm tot}^{2}\rangle$ and an estimate $\overline{\langle Z_{\rm tot}^{2}\rangle}$ from either fidelity extrapolation (blue) or bond dimension extrapolation (green) plotted against fidelity for $\Delta\theta=2\pi/9$ with varying system size. Fidelity here means the fidelity of the highest bond-dimension point used for the extrapolation. The main plots show all simulation results (individual dots) for 1 up to 40 Trotter steps together with the average relative errors (straight lines) obtained by binning the data points by fidelity using bins of width 0.05. The shaded areas signify a confidence interval of $\pm$ one standard deviation within each fidelity bin. The embedded plots show the mean values of the absolute relative errors against fidelity.
@@
By looking at the dataset of all simulations run for $\Delta\theta=2\pi/9$ on system sizes up to $6\times 6$, we find that the observation from Fig. A11 holds more generally: extrapolating fidelity $F_{\text{sim}}\rightarrow 1$ provides more reliable extrapolations than $1/\chi\rightarrow 0$ does  [^4] The fact that fidelity allows for more accurate extrapolations than bond dimension does make intuitive sense: the fidelity contains information on the accuracy of the simulation, whereas the bond dimension merely specifies the amount of classical resources used.. This can be seen in Fig. A12, which shows the relative error obtained by either fidelity or bond dimension extrapolation compared to the exact value of $\langle Z_{\rm tot}^{2}\rangle$ obtained via a statevector simulation at $\Delta\theta=2\pi/9$ for varying system sizes. Every single point in the scatter plots of Fig. A12 represents a set of simulations for a list of bond dimensions $\chi_{1},\chi_{2},\ldots,\chi_{m}$, where the x-value is the fidelity obtained by the (highest-bond-dimension) $\chi_{m}$ simulation, and the y-value is the relative error $\left(\langle Z_{\rm tot}^{2}\rangle-\overline{\langle Z_{\rm tot}^{2}\rangle}\right)/\left|\langle Z_{\rm tot}^{2}\rangle\right|$ between the exact value $\langle Z_{\rm tot}^{2}\rangle$ and an estimate $\overline{\langle Z_{\rm tot}^{2}\rangle}$ obtained using either $1/\chi$- (green) or fidelity- (blue) extrapolation via a straight line fit through the $\chi_{m-2}$, $\chi_{m-1}$ and $\chi_{m}$ points. The solid lines in Fig. A12 represent the mean of the relative errors obtained by binning the data points according to their fidelities, using a bin width of 0.05. All points in the shaded areas are within one standard deviation of the mean error. The embedded plots in Fig. A12 show the means of the absolute relative errors as a function of fidelity.
@@
From the main plots in Fig. A12 we can see that extrapolations in one over bond dimension tend to overestimate the value of $\langle Z_{\rm tot}^{2}\rangle$, and extrapolations in fidelity tend to underestimate its value. For the larger system sizes ($5\times 6$ and $6\times 6$), the mean relative error for fidelity extrapolations is always positive. From the embedded plots in Fig. A12 we can conclude that for simulations for which the highest fidelity data point used for the straight line fit has a fidelity of more than $\sim 0.65$ (i.e. to the right of where the blue and green lines cross), fidelity extrapolation is on average more accurate than bond-dimension extrapolation. Moreover, we find that as the system size grows, this cross-over point moves leftward, suggesting that fidelity extrapolation becomes more accurate with increasing system size compared to bond dimension extrapolation. Left of where the blue and green lines cross, bond-dimension extrapolation actually performs better than fidelity extrapolation. However, in this regime, the extrapolations themselves are inaccurate (as can be seen from the relative errors). Given the fidelities obtained for the quench simulations shown in Figs. (A14,A16), we believe fidelity extrapolation likely performs best for the quenches of interest, and so have obtained all extrapolated estimates in this manner.
@@
Using data obtained from simulations on smaller systems, we can add a (heuristic) notion of accuracy to the extrapolated values. In Fig. A12 we observe a clear relationship between the fidelity (of the highest bond dimension simulation performed) and the relative error on $\langle Z_{\rm tot}^{2}\rangle$. Focusing on the largest $6\times 6$ simulations, which are closest to and therefore hopefully representative of the $N=56$ qubit simulations, we obtain an average (one-sided) relative error on the extrapolation as a function of fidelity (given by the blue line in the $6\times 6$ plot of Fig. A12), together with a standard deviation on this same error (the shaded blue region in the $6\times 6$ plot of Fig. A12 lies within one $\sigma$ of the average error).
The shaded purple region in Fig. A16 signifies a one-$\sigma$ confidence interval around the extrapolated value of $\langle Z_{\rm tot}^{2}\rangle$. Because the average error (inferred from Fig. A12) is always positive, the confidence interval lies strictly above the extrapolated value. The shaded purple region is bounded from above by the extrapolated value plus average error plus $\sigma$, and from below by the smallest of (i) the extrapolated value plus average error minus $\sigma$ and (ii) the extrapolated value itself. The reason for not simply using (i) as a lower bound for the shaded region is that it is often larger than the extrapolated value itself. Given that the shaded purple area shows a confidence region of the extrapolated value, we have extended it so that it borders the extrapolated value itself. The confidence region is not shown beyond $s=10$, because the estimated errors become too large (at the low fidelities obtained) for the extrapolated values to be reliable.
The same extrapolated value and corresponding one-$\sigma$ confidence region are also shown in Fig. 3. Comparing to the zero-noise extrapolated data from H2 displayed in Fig. 3, we see that, indeed, the zero-noise extrapolated data points consistently lie within the shaded purple confidence region. The confidence region is not displayed in Fig. A14 because it is too small to be clearly visible given the high fidelity of the simulations (here, the width of the one-$\sigma$ confidence region is bounded by $0.01$ for all Trotter times).
@@
Figure A17: The expectation value of $\langle Z_{\textrm{tot}}^{2}\rangle$ from the OBC-PEPS simulation using the FU algorithm and the BP algorithm. In addition, we compute the total truncation error as the accumulated truncation error after each gate applications in the second row. In the third row, we compute the rescaled Rényi-$2$ entanglement entropy $S^{[2]}/L_{x}$ at half cut using estimation constructed from the boundary MPO. The H2 data shown here is with ZNE + ZNR.
@@
A5.1 PEPS simulation with Full Update algorithm

### v2

@@ II Floquet prethermalization in the 2D Ising model
The unitary $\mathscr{U}$ of a single Trotter layer can be realized with a quantum circuit consisting of two layers of single-qubit gates [each layer comprised of a $U_{X}(hdt)$ gate on every qubit, where $U_{X}(\varphi)=\exp(-i\varphi X/2)$] enclosing a fixed number of layers of two-qubit gates that depends on the lattice connectivity (Fig. 1b). For the simulations in this work we use a 2D rectangular lattice with periodic boundary conditions, and therefore each Trotter step consists of four layers of ($N/2$) two-qubit gates $U_{ZZ}(2Jdt)$, where $U_{ZZ}(\varphi)=\exp(-i\varphi Z\otimes Z/2)$. The $U_{ZZ}(\varphi)$ gates are implemented natively on the H2 quantum computer (i.e., they are not further decomposed into perfect entanglers).
@@ III Quantum and classical simulations
Quantum and classical simulation results for the low temperature ($\Delta\theta=0$) quench are shown in Fig. 3a. The semi-transparent solid lines are obtained using MPS-based quantum circuit simulations [27] employing a method similar to that described in Ref. [28] (see Sec. A4 for further details). We carried out these simulations up to a maximum MPS bond dimension of $\chi=4000$, for which MPS-based truncation errors were sufficiently small that we could reliably extrapolate the largest-bond dimension results to the limit of unity state fidelity (purple line, zero-truncation extrapolation or ZTE, Sec. A4). We also show raw (completely unprocessed) data from the H2 quantum computer. The dominant source of error in these simulations is imperfections in the two qubit gates, which have an average fidelity of $99.94(1)\%$ (see Sec. A2.2). We also obtained data for noise amplified (NA) circuits, in which two-qubit Pauli operators are stochastically inserted (based on a learned noise channel of our two-qubit gates) in order to artificially reduce the quality of the quantum computation [12]. From the raw and NA data, the quantum data is extrapolated to the limit of zero noise using zero-noise extrapolation (ZNE) [29]. Details of the ZNE methods, along with a novel strategy to remove the impact of detectable leakage errors without post-selection (zero-noise regression, or ZNR), are presented in the appendices. The good agreement of the processed data with the extrapolated MPS results gives us confidence in the error-mitigation strategies in a regime where they have considerable error to remove.
@@
Figure 3c shows the same error-mitigated quantum data for the intermediate temperature quench compared to exact diagonalization results for smaller systems (gray curves), along with further approximate classical methods for the $7\times 8$ lattice, providing a comprehensive comparison of the quantum data with a variety of classical approaches at the frontier of numerical simulations. We explored two different classes of methods: Schrödinger picture methods that evolve a compressed representation of the wave function, and Heisenberg picture methods that evolve a compressed representation of the observable $Z_{\rm tot}^{2}$. For the former, we attempted to evolve the wave function using a projected entangled pair state (PEPS) ansatz and a variety of compression schemes (including full update and belief propagation). We found that evolving a PEPS with periodic-boundary conditions using belief propagation (BP-PEPS), and then computing $\langle Z_{\rm tot}^{2}\rangle$ by sampling, produced accurate observables to the longest times, though we were still not able to obtain converged results beyond about 8 Trotter steps (mainly due to GPU memory limitations, see Sec. A5.2). Simulations in the Heisenberg picture were carried out using both projected entangled pair operator (PEPO, see Sec. A5.3) and sparse-Pauli methods [33, 34, 35, 36, 37, 38] (see Sec. A6 comparing coefficient and Pauli weight truncations). In general, we found PEPO evolution to be less accurate than PEPS simulations with comparable resources. It would be interesting to combine Schrödinger and Heisenberg techniques [33, 39], though the limitations of each independently for this quench suggests that a hybrid approach would likely face challenges in substantially extending the temporal range. Sparse Pauli dynamics simulations retained about 4 billion Pauli operators and used roughly $1{\rm TB}$ of memory, and while they are in principle systematically improvable with more memory they could not be converged with available resources. We also performed neural network quantum state (NQS) simulations using the convolutional neural network (CNN) architecture similar to Ref. [40]. The NQS is time evolved by time-dependent variational Monte Carlo (tVMC) [41] using the NetKet library [42]. By systematically increasing the number of parameters of the ansatz, we obtained good results for smaller systems, but we were not able to obtain quantitatively accurate results except at very short times for a $7\times 8$ lattice (detailed results are presented in Appendix A7). NQS-based methods are a rapidly developing field of study, and it would be interesting to see whether any of the many possible variations can be brought in agreement with the quantum data out to late times.
@@ IV Emergent hydrodynamics
Figure 4: Emergent hydrodynamics. Hydrodynamic relaxation of the spin-exchange contribution to energy density for a quench in an $14\times 4$ strip with an inhomogeneous initial state (described in text). (a) Evolution of the y-averaged exchange energy density profiles, $\mathcal{E}(x)$. (b) Dashed lines show a fit of $\log|\tilde{\mathcal{E}}(q)|^{2}$ to $\log(ae^{-\Gamma_{q}s}+c\delta_{q,0})$. Fits are restricted to long-wavelengths ($q=2\pi n/L_{x}$ for $n=0,1,2,3$), and exclude data points with poor signal-to-shot-noise ratio (unfilled symbols). The extracted decay rates are consistent with the expected diffusive scaling of the heat equation, $\Gamma_{q}\approx\mathcal{D}q^{2}$ (inset) with diffusion constant $\mathcal{D}=0.38(5)$.
An essential aspect of prethermalization in Floquet systems is the emergence of an approximately conserved energy, whose violation only becomes apparent at exponentially late times, $t\gtrsim\tau_{H}$. In chaotic (pre)thermalizing systems, local observables that overlap with the conserved energy density are expected to exhibit a short-time transient relaxation to a local thermal quasi-equilibrium followed by a slow, hydrodynamic relaxation of long-wavelength energy inhomogeneities at late times. To probe the expected emergence of hydrodynamics, we simulate a quench with an inhomogeneous initial condition in which a thin vertical domain of down (|$\ket{1}$) spins is embedded in an otherwise uniform background, corresponding to Eq. (2) with site dependent $\theta_{j}=\pi$ if $x_{j}=L_{x}/2$ or otherwise $\theta_{j}=\theta_{\rm min}+2\pi/9$ ($x_{j}$ being the $x$-coordinate of qubit $j$). To create an appropriate separation of time-scales between the local thermalization time ($t_{0}\gtrsim 5$ time steps for these parameters), and late-time hydrodynamics, we consider an elongated rectangular geometry with dimensions $(L_{x},L_{y})=(14,4)$. We measure on the quantum processor the local density of $H_{ZZ}$: $\mathcal{E}_{i}=\frac{J}{4}\sum_{j\in\langle i\rangle}\langle Z_{i}Z_{j}\rangle$, average this quantity over the $y$ coordinate, $\mathcal{E}(x)=\frac{1}{L_{y}}\sum_{i:x_{i}=x}\mathcal{E}_{i}$ (see Fig. 4a), and compute its 1D Fourier transform ~$\tilde{\mathcal{E}}(q)=\sum_{x}\frac{e^{iqx}}{\sqrt{L_{x}}}\mathcal{E}(x)$ (see Fig. 4b). The $q=0$ mode quickly saturates to a steady-state value, indicating its overlap with an emergent conserved energy. The smallest few non-zero ($q_{n}=2\pi n/L_{x}$, $n=1,2,3$) wavevectors relax with characteristic rates $\Gamma_{q}$, consistent with the expected behavior for the heat equation, $\Gamma_{q}\approx\mathcal{D}q^{2}$ (valid asymptotically for long wavelengths, $|{q}|\ll 1$), with diffusion constant $\mathcal{D}=0.38(5)$. Components with larger $q$ decay rapidly within the initial local thermalization and do not survive into the hydrodynamic regime.
@@ Acknowledgments
We acknowledge the entire Quantinuum team for their many contributions towards successful operation of the H2 quantum computer, and Honeywell for fabricating the trap used in this experiment. We thank Konstantinos Meichanetzidis, Chris Langer, Grahame Vittorini, Stephen Erickson, James Hostetter, and David Liefer for helpful comments on the manuscript and thank Dan Mills for useful discussions. S.A.M. contributed to the construction of the experimental apparatus while at Quantinuum. Fermioniq was supported by the Dutch National Growth Fund (NGF), as part of the Quantum Delta NL Programme. W.K., M.K., and F.P. acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy–EXC–2111–390814868, TRR 360 – 492547816 and DFG grants No. KN1254/1-2, KN1254/2-1, FOR 5522 (project-id 499180199), the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 851161), the European Union (grant agreement No 101169765), as well as the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus. E.G. acknowledges funding by the Bavarian Ministry of Economic Affairs, Regional Development and Energy (StMWi) under project Bench-QC (DIK0425/01). M.S.R. acknowledges funding from the 2024 Google PhD Fellowship, the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (grant number 205602), and the the Swiss National Science Foundation (grant number 200021-219329). SPD simulation work at Caltech was supported by the US Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. TN-BP algorithm development at Caltech was supported by the US Department of Energy, Office of Science, Basic Energy Sciences and Advanced Scientific Computing Research, Quantum Utility through Advanced Computational Quantum Algorithms.
@@
which performs a rotation by $0\leq\theta\leq\pi$ about a Bloch vector in the $xy$ plane specified by the angle $0\leq\phi<2\pi$. (The $U_{X}(\varphi)$ gate in our Trotter circuit, shown in Fig. 1, is implemented as $U_{X}(\varphi)=U_{1q}(\varphi,0)$.) An $X$ gate used in DD can be implemented in two physically inequivalent ways: $U_{1q}(\pi,0)=-iX$ and $U_{1q}(\pi,\pi)=+iX$. While these two gates differ by a logically-irrelevant global phase, coherent errors on the two gates due to systematic under- or over-rotations are not logically equivalent. During DD, we alternate the phase of each $X$ gate on each qubit between $U_{1q}(\pi,0)$ and $U_{1q}(\pi,\pi)$ to approximately cancel such errors.
@@ A Pauli error channel is given by
In this section we present the classical simulation results of the digitized TFIM obtained with 1D matrix-product-state (MPS) [57, 58, 59] methods. The results are accompanied by a quantitative analysis of the extrapolation method used, from which we conclude that, for the observable of interest [$Z_{\rm tot}^{2}$ from Eq. (4)], extrapolating in simulation fidelity is more reliable than extrapolating in bond dimension – an observation that likely holds more generally beyond the scope of this work. Furthermore, based on comparisons to exact simulations for smaller systems, we include a heuristic confidence interval on the extrapolations. The zero-noise extrapolated expectation values obtained from H2 are found to lie within this confidence interval at the early times for which it can be reliably established (i.e., before the MPS fidelity becomes so low that the extrapolations are completely unpredictable, up to $s\approx 8$ for the intermediate quench).
@@
Figure A12: Relative error $\left(\langle Z_{\rm tot}^{2}\rangle-\overline{\langle Z_{\rm tot}^{2}\rangle}\right)/\left|\langle Z_{\rm tot}^{2}\rangle\right|$ between the exact value $\langle Z_{\rm tot}^{2}\rangle$ and an estimate $\overline{\langle Z_{\rm tot}^{2}\rangle}$ from either fidelity extrapolation (blue) or bond dimension extrapolation (green) plotted against fidelity for $\Delta\theta=2\pi/9$ with varying system size. Fidelity here means the fidelity of the highest bond-dimension point used for the extrapolation. The main plots show all simulation results (individual dots) for 1 up to 40 Trotter steps together with the average relative errors (straight lines) obtained by binning the data points by fidelity using bins of width 0.05. The shaded areas signify a 5th - 95th percentile window within each fidelity bin. The embedded plots show the mean values of the absolute relative errors against fidelity.
@@
By looking at the dataset of all simulations run for $\Delta\theta=2\pi/9$ on system sizes up to $6\times 6$, we find that the observation from Fig. A11 holds more generally: extrapolating fidelity $F_{\text{sim}}\rightarrow 1$ provides more reliable extrapolations than $1/\chi\rightarrow 0$ does  [^4] The fact that fidelity allows for more accurate extrapolations than bond dimension does make intuitive sense: the fidelity contains information on the accuracy of the simulation, whereas the bond dimension merely specifies the amount of classical resources used.. This can be seen in Fig. A12, which shows the relative error obtained by either fidelity or bond dimension extrapolation compared to the exact value of $\langle Z_{\rm tot}^{2}\rangle$ obtained via a statevector simulation at $\Delta\theta=2\pi/9$ for varying system sizes. Every single point in the scatter plots of Fig. A12 represents a set of simulations for a list of bond dimensions $\chi_{1},\chi_{2},\ldots,\chi_{m}$, where the x-value is the fidelity obtained by the (highest-bond-dimension) $\chi_{m}$ simulation, and the y-value is the relative error $\left(\langle Z_{\rm tot}^{2}\rangle-\overline{\langle Z_{\rm tot}^{2}\rangle}\right)/\left|\langle Z_{\rm tot}^{2}\rangle\right|$ between the exact value $\langle Z_{\rm tot}^{2}\rangle$ and an estimate $\overline{\langle Z_{\rm tot}^{2}\rangle}$ obtained using either $1/\chi$- (green) or fidelity- (blue) extrapolation via a straight line fit through the $\chi_{m-2}$, $\chi_{m-1}$ and $\chi_{m}$ points. The solid lines in Fig. A12 represent the mean of the relative errors obtained by binning the data points according to their fidelities, using a bin width of 0.05. All points in the shaded areas lie within the 5th - 95th percentile window of the errors obtained within each bin. The embedded plots in Fig. A12 show the means of the absolute relative errors as a function of fidelity.
@@
From the main plots in Fig. A12 we can see that extrapolations in one over bond dimension tend to overestimate the value of $\langle Z_{\rm tot}^{2}\rangle$, and extrapolations in fidelity tend to underestimate its value. From the embedded plots in Fig. A12 we can conclude that for simulations for which the highest fidelity data point used for the straight line fit has a fidelity of more than $\sim 0.65$ (i.e. to the right of where the blue and green lines cross), fidelity extrapolation is on average more accurate than bond-dimension extrapolation. Moreover, we find that as the system size grows, this cross-over point moves leftward, suggesting that fidelity extrapolation becomes more accurate with increasing system size compared to bond dimension extrapolation. Left of where the blue and green lines cross, bond-dimension extrapolation actually performs better than fidelity extrapolation. However, in this regime, the extrapolations themselves are inaccurate (as can be seen from the relative errors). Given the fidelities obtained for the quench simulations shown in Figs. (A14,A16), we believe fidelity extrapolation likely performs best for the quenches of interest, and so have obtained all extrapolated estimates in this manner.
@@
Using data obtained from simulations on smaller systems, we can add a (heuristic) notion of accuracy to the extrapolated values. In Fig. A12 we observe a clear relationship between the fidelity (of the highest bond dimension simulation performed) and the relative error on $\langle Z_{\rm tot}^{2}\rangle$. Focusing on the largest $6\times 6$ simulations, which are closest to and therefore hopefully representative of the $N=56$ qubit simulations, we obtain an average relative error on the extrapolation as a function of fidelity (given by the blue line in the $6\times 6$ plot of Fig. A12), together with a 5th - 95th percentile window around this same error (visualized by the shaded blue region in the $6\times 6$ plot of Fig. A12).
The shaded purple region in Fig. A16 signifies a confidence interval around the extrapolated value of $\langle Z_{\rm tot}^{2}\rangle$ based on the 5th - 95th percentile window of the relative errors of the $6\times 6$ simulations. The confidence region is not shown beyond $s=10$, because once the estimated errors become too large and unpredictable (at the low fidelities obtained) the confidence interval itself becomes less meaningful. The same extrapolated value and corresponding confidence region are also shown in Fig. 3. Comparing to the zero-noise extrapolated data from H2 displayed in Fig. 3, we see that, indeed, the zero-noise extrapolated data points consistently lie within the shaded purple confidence region.
@@
Figure A17: The expectation value of $\langle Z_{\textrm{tot}}^{2}\rangle$ from the OBC-PEPS to simulate PBC states using the FU algorithm and the BP algorithm. In addition, we compute the total truncation error as the accumulated truncation error after each gate applications in the second row. In the third row, we compute the rescaled Rényi-$2$ entanglement entropy $S^{[2]}/L_{x}$ at half cut using an estimation constructed from the boundary MPO. The H2 data shown here is with ZNE + ZNR.
@@
A5.1 PEPS simulation with the Full Update algorithm